# Strongly Interacting Dark Matter admixed Neutron Stars

### v2

@@ 2.1 Ordinary matter
For ordinary matter we work with the equations of state from Kurkela:2014vha . This work uses piecewise polytropes, to interpolate between a low density regime described by nuclear chiral perturbation theory (NChPT) Hebeler_2013 and a high density regime described by perturbative QCD Kurkela:2009gj . Equations of state that do not meet the observational constraints from NICER data on the mass, radius and the tidal deformability are omitted.
For this study, we use 3 limiting equations (EoS I, II and III) of state from Kurkela:2014vha . EoS I is the softest of the three, barely reaching the two solar mass limits. EoS II yields a maximal maximum mass. EoS III is the stiffest and while it might be too stiff to describe an ordinary neutron star, we still include it in our analysis as the addition of dark matter to a neutron star can result in smaller masses and radii. Therefore it might be able to meet observational constraints when considering mixed neutron stars. For the same reason we will employ central pressures larger than the ones stated in Kurkela:2014vha that result in the maximal mass. We will see that the addition of dark matter allows for stable solutions at larger central pressures Kain:2021hpk ; Valdez-Alvarado:2012rct . Figure 1 shows the different equations of state alongside the equation of state used to describe dark matter.
We extend the equations of state below the NChPT datapoints in Kurkela:2014vha by interpolating from zero using another polytrope with adiabatic index $\Gamma=5/3$. This corresponds to a free fermi gas and coincides with the equation of state $n\propto(\mu^{2}-m_{n}^{2})^{3/2}$ for low densities. We do not include a specified equation of state for the crust which is sufficient for the scope of this work, as the crust would hardly affect the radius and the effect on the mass would be negligible.
@@ 2.2 G2-QCD
G2-QCD is a QCD-like theory with the gauge group SU(3) replaced by the exceptional Lie group G2  Holland:2003jy ; Pepe:2006er . We consider the situation with one flavor of fermions in the fundamental representation Maas:2012wr ; Wellegehausen:2013cya , called dark quarks in the following. G2-QCD has, due to its group structure, a number of extraordinary features. Most relevant here is that chiral symmetry breaking occurs already in the one-flavor theory, and that dark baryons can be built with any number of dark quarks Kogut:2000ek ; Holland:2003jy ; Wellegehausen:2013cya . It can also be simulated at all densities using first-principle lattice methods Maas:2012wr using standard algorithms Gattringer:2010zz ; Wellegehausen:2011sc .
@@
A detailed discussion of its spectrum and other properties in the vacuum can be found in  Holland:2003jy ; Pepe:2006er ; Wellegehausen:2013cya . The most relevant features for our purpose here is that the lightest absolutely stable fermionic state is supposedly a three dark-quark state. In addition, there can exist a one dark-quark hadron  Holland:2003jy , but it is likely heavier for the dark quark masses employed here, because it is a dark-quark-dark-gluon hybrid state Hajizadeh:2017jsw ; Wellegehausen:2013cya . It is thus supposedly unstable against decay into the three dark-quark state and dark mesons. Note that in G2-QCD dark quarks and dark antiquarks, as e. g. also in Sp(4)-QCD and SU(2)-QCD, cannot be distinguished Kogut:2000ek , and thus only Grassmann parity is conserved. In fact, the dark baryonic chemical potential is a Weyl-flavor potential, but serves the same purpose as the baryonic chemical potential in QCD Maas:2012wr ; Wellegehausen:2013cya . As a consequence, dark mesons cannot be distinguished from dark glueballs in the one-flavor theory, and thus are only stable if the theory remains in isolation. Hence, we identify the three dark-quark state as the dark matter candidate of our theory. We therefore use its mass also for scale setting.
@@
Details on the equation of state, its determination and properties, can be found in  Wellegehausen:2013cya ; Maas:2012wr ; Hajizadeh:2017jsw . Here, we concentrate on the equation of state at zero temperature, which has been discussed in Wellegehausen:2013cya ; Hajizadeh:2017jsw . Of course, the lattice data is only available at discrete values of the dark baryonic chemical potential. The interpolation between data points is, as for the ordinary matter equation of state, performed using piecewise polytropes to ensure thermodynamic consistency.
@@
Just like for the ordinary matter equation of state we use choose $\Gamma=5/3$ for the polytrope describing the low density behavior. This coincides with $n=c(\mu^{2}-m_{C}^{2})^{3/2}$ for low densities, where $m_{C}$ is the mass of the dark matter candidate and c is a free parameter as was considered in Hajizadeh:2017jsw . For the light ensemble and at high densities, we extrapolate the data using the fit form $n(\mu)=n_{s}/(\exp(a-b\mu)+1)$ fitted to the last few data points Wellegehausen:2013cya ; Hajizadeh:2017jsw . Herein $n_{s}$ is the lattice saturation density, i. e. the situation where Pauli blocking obstructs larger densities. This is a pure lattice artifact but the densities required in the following, as has already been observed previously Hajizadeh:2017jsw , are small enough for this to be not a relevant issue. In the heavy ensemble, the extrapolation region is never probed. The interpolation and extrapolation captures the main physical features of the lattice results while providing access to a larger range in the density. Note that, in contrast to the low-density case, the high-density fit function has not been chosen on physical grounds, but for its ability to describe the lattice data very accurately.
The equation of state has been obtained for two different values of the bare lattice parameters [^2] The only free parameter in the theory for the dark matter is the bare fermion mass., which will result in different mass spectra. The lightest particle in the theory is the pseudo-scalar dark pion which is unstable to potential decays to the standard model via messengers. The dark matter candidate is the lightest fermionic particle. In the two resulting ensembles the mass of the dark pion is 0.25 $m_{c}$ and 0.33 $m_{c}$ respectively. The original choice of the lattice simulation parameters were for a different reasons than dark matter Maas:2012wr , and thus do not apply here. It would require a major effort and incur computational costs of the order of a few tens of millions of core hours to provide further values  Wellegehausen:2013cya . This would only be justified by concrete observational hints for a particular mass value. However, for our exploratory investigations, the precise value does not matter too much. As will be seen, the impact is only a limited quantitative one, and the change between the ordinary matter EoSs has already a more substantial effect.
The resulting dark matter equation of states are also shown in figure 1.
@@ 3 Building a two-component neutron star
In this section we will describe the framework used to build a neutron star from the two-fluid TOV-equation Tolos:2015qra ; Hajkarim:2024ecp . For comparison, we depict the different involved equations of state in figure 1.
@@ 3.1 Two-fluid TOV
The TOV-equation is a a system of coupled differential equations describing the pressure gradient $dp/dr$ and mass gradient $dm/dr$ derived from the conservation of the energy stress tensor and is consistent with general relativistic effects. Assuming a static, spherically symmetric object in hydrostatic equilibrium, one obtains the ordinary TOV-equation consisting of the pressure and a mass gradient. While usually used for determining ordinary matter properties, a set of two-fluid TOV-equations containing an exotic matter component can be achieved by neglecting the interactions between ordinary and exotic matter components. The following set of equations are the two-fluid TOV-equations written in a dimensionless form Narain:2006kx ; Tolos:2015qra . In particular, after switching to natural units, we rescale all dimensionful quantities in units of the mass of the dark matter candidate $m_{C}$, yielding
@@
where the subscript $O$ refers to the ordinary matter fluid [^3] Sometimes this is also called baryonic matter. We decide to use ordinary ($O$) as our dark matter is also baryonic. while $D$ stands for dark matter fluid. $\varepsilon$ is the energy density. In the following, lowercase $m$ and $r$ will refer to the mass and radius during the integration while capital letters $M$ and $R$ refer to neutron star properties. Finally, the metric function $d\nu/dr$ is defined for both fluids as
@@
The coupled two-fluid TOV equations can be solved using the equations of state of the two fluids. Iterating over the potentially nonidentical central pressures, one obtains the neutron star properties for each combination of central pressures ($p_{0,O}$ and $p_{0,D}$) by integrating the equations until the pressure vanishes. The radius at which the ordinary/dark matter pressure drops to zero within numerical precision is labeled ROR{O}/RDR{D}. $M_{O}$$M_{D}$ are the integrated masses of the two fluids respectively and $M_{tot}=M_{O}+M_{D}$. The solutions can result in a neutron star with a dark core ($r_{O}>r_{D}$) or a dark halo ($r_{O}<r_{D}$), which potentially also affects the gravitational wave signal of an inspiral of two neutron stars Ivanytskyi:2019wxd .
Dimensional analysis shows that the mass and the radius scale like $1/m_{N}^{2}$. Given an equation of state that can be trivially scaled by the mass of its constituents $m_{const}$, as is the case with our dark matter equation of state, halving the mass of the constituents would result in a star with four times the mass and four times the radius. To some extent, this scaling argument is still present in the two fluid case. A lighter dark matter candidate mass will result in a stronger impact of dark matter on the star. In the limiting cases where the central pressure of one of either the dark or ordinary matter is negligible ($p_{0,D}\gg p_{0,O}$ or $p_{0,D}\ll p_{0,O}$) this scaling argument is perfectly recovered as very heavy dark matter candidates will have no impact on the observables while a light dark matter candidate will result in a compact object dominated by dark matter that can eventually not be interpreted as a neutron star anymore.
@@ 3.2 Stability
Not every solution of the TOV equations yields a stable neutron star. Small perturbations in the metric field should settle back to the original solution. By solving a Sturm-Liouville problem one can identify the stable solutions by demanding the following criterion Hippert:2022snq ; Pitz:2024xvh ; Barbat:2024yvi
@@
where, K is a constant and $\Gamma$ depends on the polytropic index n like $\Gamma=\frac{\text{n}+1}{\text{n}}$ Sagert:2005fw .
@@ 3.3 Tidal Deformability
When two compact objects orbit each other, they get deformed by the gravitational tidal field of one another. The tidal deformability $\Lambda$ is a quantity that describes the proportionality between the tidal field and the induced quadrupole deformation. We will work with the unitless version of the tidal deformability that is connected to the dimensionful tidal deformability $\lambda$ and the dimensionless second Love number $k_{2}$ via
@@
Here, we have introduced the compactness $C=M_{tot}/R$ calculated with the total mass $M_{tot}=M_{O}+M_{D}$ and the maximal radius $R=\max(R_{O},R_{D})$.
Calculating the tidal deformability (eqn. 6) requires the second Love number, which can be calculated from the neutron star compactness as
@@
with $m=m_{O}+m_{D}$ and $c_{s,O/D}^{2}=dP_{O/D}/d\varepsilon_{O/D}$ is the speed of sound Barbat:2024yvi .
In contrast to black holes, neutron stars have a non-vanishing tidal deformability. As a result, the gravitational wave signal of a neutron star binary differs from the one produced by a black hole merger by a (potentially time-dependent) phase in the late stages of the inspiral. To leading order in the tidal deformabilities of the two neutron stars, this phase can be determined by the parameter $\tilde{\Lambda}$ as LIGO_GW170817
@@
where $\Lambda_{1}$ and $\Lambda_{2}$ are the tidal deformabilities and $M_{1}$ and $M_{2}$ are the total masses of the two neutron stars. Thus, deviations in this quantity compared to an ordinary matter neutron star give a measure of the impact a dark matter component will have on the gravitational wave signal of two compact stellar objects. As the observed phase difference cannot separate both $\Lambda_{i}$, we choose conservatively the extreme cases that either one entirely dominates for the limits in $\Lambda$ below.
The applicability of gravitational wave analysis beyond black-hole signals to more complex and exotic systems such as the mixed stars studied here is not thoroughly understood Cardoso:2016rao ; Cardoso:2017cqb ; DiGiovanni:2022mkn ; Bauswein:2020kor . As the nature of this work is exploratory, we will discuss the results of the tidal deformability for all our results for illustrative purposes only.

### v3

@@ 2.1 Ordinary matter
For ordinary matter we work with the EoS from Kurkela et al. (2014). This work uses piecewise polytropes, to interpolate between a low density regime described by nuclear chiral perturbation theory (NChPT) Hebeler et al. (2013) and a high density regime described by perturbative QCD Kurkela et al. (2010). Perturbative QCD is only valid at densities much larger than those found in neutron stars, but asymptotically will describe the EoS. Thus, following Kurkela et al. (2014), an interpolation between the two regimes should be possible within limits like the speed of sound $c_{s}$ satisfying $c_{s}<1$.
For this study, we use 3 limiting equations of state (EoS I, II and III) from Kurkela et al. (2014). EoS I is the softest of the three, barely reaching the two solar mass limits. EoS II yields a maximal maximum mass. EoS III is the stiffest and while it might be too stiff to describe an ordinary neutron star, we still include it in our analysis as the addition of dark matter to a neutron star can result in smaller masses and radii. Therefore it might be able to meet observational constraints when considering mixed neutron stars. For the same reason we will employ central pressures larger than the ones stated in Kurkela et al. (2014) that result in the maximal mass. We will see that the addition of dark matter allows for stable solutions at larger central pressures Kain (2021); Valdez-Alvarado et al. (2013). Figure˜1 shows the pressure and the speed of sound versus the energy density for the three ordinary matter and two dark matter equations of state employed.
We extend the EoS below the NChPT data points in Kurkela et al. (2014) by interpolating from zero using another polytrope with adiabatic index $\Gamma=5/3$. This corresponds to a free fermi gas and coincides with the equation of state $n\propto(\mu^{2}-m_{n}^{2})^{3/2}$ for low densities. We do not include a specific EoS for the crust, which would affect the radius. This should be considered a systematic uncertainty, but likely a subleading one for the present purpose.
@@ 2.2 G2-QCD
G2-QCD is a QCD-like theory with the gauge group SU(3) replaced by the exceptional Lie group G2  Holland et al. (2003); Pepe and Wiese (2007). We consider the situation with one flavor of fermions in the fundamental representation Maas et al. (2012); Wellegehausen et al. (2014), called dark quarks in the following. G2-QCD has, due to its group structure, a number of extraordinary features. Most relevant here is that chiral symmetry breaking occurs already in the one-flavor theory, and that dark baryons can be built with any number of dark quarks Kogut et al. (2000); Holland et al. (2003); Wellegehausen et al. (2014). It can also be simulated at all densities using first-principle lattice methods Maas et al. (2012) using standard algorithms Gattringer and Lang (2010); Wellegehausen et al. (2011). This will necessarily capture all self-interaction effects on the EoS completely.
@@
A detailed discussion of its spectrum and other properties in the vacuum can be found in  Holland et al. (2003); Pepe and Wiese (2007); Wellegehausen et al. (2014). The most relevant features for our purpose here is that the lightest absolutely stable fermionic state is supposedly a three dark-quark state. In addition, there can exist a one dark-quark hadron  Holland et al. (2003), but it is likely heavier for the dark quark masses employed here, because it is a dark-quark-dark-gluon hybrid state Hajizadeh and Maas (2017); Wellegehausen et al. (2014). It is thus supposedly unstable against decay into the three dark-quark state and dark mesons. Note that in G2-QCD dark quarks and dark antiquarks, as e. g. also in Sp(4)-QCD and SU(2)-QCD, cannot be distinguished Kogut et al. (2000), and thus only Grassmann parity is conserved. In fact, the dark baryonic chemical potential is a Weyl-flavor potential, but serves the same purpose as the baryonic chemical potential in QCD Maas et al. (2012); Wellegehausen et al. (2014). As a consequence, dark mesons cannot be distinguished from dark glueballs in the one-flavor theory, and thus are only stable if the theory remains in isolation. Hence, we identify the three dark-quark state as the dark matter candidate of our theory. We therefore use its mass also for scale setting. At the current time, low-energy features beyond the spectrum in Wellegehausen et al. (2014) have not been obtained. In particular, establishing a low-energy effective theory in the spirit of Kulkarni et al. (2023) has not (yet) been done, and would require a major effort. However, the low-energy effective theory is never explicitly needed here, and all non-trivial effects on the EoS are automatically cpatured by the lattice simulations.
@@
Details on the EoS, its determination and properties, can be found in  Wellegehausen et al. (2014); Maas et al. (2012); Hajizadeh and Maas (2017). Here, we concentrate on the EoS at zero temperature, which has been discussed in Wellegehausen et al. (2014); Hajizadeh and Maas (2017). Of course, the lattice data is only available at discrete values of the dark baryonic chemical potential. The interpolation between data points is, as for the ordinary matter equation of state, performed using piecewise polytropes to ensure thermodynamic consistency.
@@
Just like for the ordinary matter EoS we use choose $\Gamma=5/3$ for the polytrope describing the low density behavior. This coincides with $n=c(\mu^{2}-m_{C}^{2})^{3/2}$ for low densities, where $m_{C}$ is the mass of the dark matter candidate and c is a free parameter as was considered in Hajizadeh and Maas (2017). For the light ensemble and at high densities, we extrapolate the data using the fit form $n(\mu)=n_{s}/(\exp(a-b\mu)+1)$ fitted to the last few data points Wellegehausen et al. (2014); Hajizadeh and Maas (2017). Herein $n_{s}$ is the lattice saturation density, i. e. the situation where Pauli blocking obstructs larger densities. This is a pure lattice artifact but the densities required in the following, as has already been observed previously Hajizadeh and Maas (2017), are small enough for this to be not a relevant issue. Indeed, for the heavy ensemble below, the extrapolation region is never probed. The interpolation and extrapolation captures the main physical features of the lattice results while providing access to a larger range in the density. Note that, in contrast to the low-density case, the high-density fit function has not been chosen on physical grounds, but for its ability to describe the lattice data very accurately.
[…]
@@ 3 Building a two-component neutron star
In this section we will describe the framework used to build a neutron star from the two-fluid TOV equation Tolos et al. (2015); Hajkarim et al. (2024). For comparison, we depict the different involved EoS in Figure˜1.
@@ 3.1 Two-fluid TOV
The TOV equation is a system of coupled differential equations describing the pressure gradient $dp/dr$ and mass gradient $dm/dr$ derived from the conservation of the energy stress tensor and is consistent with general relativistic effects. Assuming a static, spherically symmetric object in hydrostatic equilibrium, one obtains the ordinary TOV equation consisting of the pressure and a mass gradient. While usually used for determining ordinary matter properties, a set of two-fluid TOV equations containing an exotic matter component can be achieved by neglecting the interactions between ordinary and exotic matter components. The following set of equations are the two-fluid TOV equations written in a dimensionless form Narain et al. (2006); Tolos et al. (2015). In particular, after switching to natural units, we rescale all dimensionful quantities in units of the mass of the dark matter candidate $m_{C}$, yielding
@@
where the subscript $O$ refers to the ordinary matter fluid [^4] Sometimes this is also called baryonic matter. We decide to use ordinary ($O$) as our dark matter is also (dark) baryonic. while $D$ stands for dark matter fluid. $\varepsilon$ is the energy density. In the following, lowercase $m$ and $r$ will refer to the mass and radius during the integration. Capital letters $M_{D/O}$ and $R_{D/O}$ refer to the neutron star mass and radius, respectively, where subscripts $D$ and $O$ denote the corresponding total dark and ordinary values for these quantities, respectively. Finally, the metric function $d\nu/dr$ is defined for both fluids as
@@
The coupled two-fluid TOV equations can be solved using the EoS of the two fluids. Iterating over the potentially nonidentical central pressures, one obtains the neutron star properties for each combination of central pressures ($p_{0,O}$ and $p_{0,D}$) by integrating the equations until the pressure vanishes. The radius at which the ordinary/dark matter pressure drops to zero within numerical precision defines $R_{O}$$R_{D}$. $M_{O}$$M_{D}$ are the resulting integrated masses of the two fluids respectively and $M_{tot}=M_{O}+M_{D}$. The solutions can result in a neutron star with a dark core ($R_{O}>R_{D}$) or a dark halo ($R_{O}<R_{D}$), which potentially also affects the gravitational wave signal of an inspiral of two neutron stars Ivanytskyi et al. (2020).
Dimensional analysis shows that the mass and the radius scale like $1/m_{N}^{2}$. Given an equation of state that can be trivially scaled by the mass of its constituents $m_{const}$, as is the case with our dark matter EoS, halving the mass of the constituents would result in a star with four times the mass and four times the radius. To some extent, this scaling argument is still present in the two fluid case. A lighter dark matter candidate mass will result in a stronger impact of dark matter on the star. In the limiting cases where the central pressure of one of either the dark or ordinary matter is negligible ($p_{0,D}\gg p_{0,O}$ or $p_{0,D}\ll p_{0,O}$) this scaling argument is perfectly recovered as very heavy dark matter candidates will have no impact on the observables while a light dark matter candidate will result in a compact object dominated by dark matter that can eventually not be interpreted as a neutron star anymore.
@@ 3.2 Stability
Not every solution of the TOV equations yields a stable neutron star. Small perturbations in the metric field should settle back to the original solution. By solving a Sturm-Liouville problem one can identify the stable solutions by demanding the following criterion Hippert et al. (2023); Pitz and Schaffner-Bielich (2024); Barbat et al. (2024)
@@
where, K is a constant and $\Gamma$ depends on the polytropic index n like $\Gamma=\frac{\text{n}+1}{\text{n}}$ Sagert et al. (2006).
@@ 3.3 Tidal Deformability
When two compact objects orbit each other, they get deformed by the gravitational tidal field of one another. The tidal deformability $\Lambda$ is a quantity that describes the proportionality between the tidal field and the induced quadrupole deformation. We will work with the dimensionless version of the tidal deformability that is connected to the dimensionful tidal deformability $\lambda$ and the dimensionless second Love number $k_{2}$ via
@@
Here, we have introduced the compactness $C=M_{tot}/R$ calculated with the total mass $M_{tot}=M_{O}+M_{D}$ and the maximal radius $R=\max(R_{O},R_{D})$. Calculating the tidal deformability requires the second Love number, which can be calculated from the neutron star compactness as
@@
with $m=m_{O}+m_{D}$ and $c_{s,O/D}^{2}=dP_{O/D}/d\varepsilon_{O/D}$ is the speed of sound Barbat et al. (2024).
In contrast to black holes, neutron stars have a non-vanishing tidal deformability. As a result, the gravitational wave signal of a neutron star binary differs from the one produced by a black hole merger by a (potentially time-dependent) phase in the late stages of the inspiral. To leading order in the tidal deformabilities of the two neutron stars, this phase can be determined by the parameter $\tilde{\Lambda}$ as Abbott et al. (2017)
@@
where $\Lambda_{1}$ and $\Lambda_{2}$ are the tidal deformabilities and $M_{1}$ and $M_{2}$ are the total masses of the two neutron stars. This definition is chosen such that $\tilde{\Lambda}=\Lambda_{1}=\Lambda_{2}$ if $M_{1}=M_{2}$. Therefore, we can use it to constrain our results at the mass extracted from the merger event. We use the low-spin prior results for which this quantity is constraint to <800 for $M_{\text{tot}}=1.38^{+0.24}_{-0.19}M_{\odot}$ from GW170817 Abbott and others (2017a, b) and <600 for $M_{\text{tot}}=1.6^{+0.27}_{-0.14}M_{\odot}$ from GW190425 Abbott and others (2020). Deviations in this quantity compared to an ordinary matter neutron star give a measure of the impact a dark matter component will have on the gravitational wave signal of two compact stellar objects.
The applicability of gravitational wave analysis beyond black-hole signals to more complex and exotic systems such as the mixed stars studied here is not fully understood Cardoso et al. (2016); Cardoso and Pani (2017); Di Giovanni et al. (2022); Bauswein et al. (2023). As the nature of this work is exploratory, we will discuss the results of the tidal deformability for all our results for illustrative purposes only.